OpenStudy (anonymous):

What is the inverse of the following equation x^2 +4 when x >0 is one to one ?

OpenStudy (anonymous):

The equation can be written as y = x^2 +4 the inverse is x = y^2 +4

OpenStudy (mertsj):

\[x=y^2+4\] \[x-4=y^2\] \[y=+\sqrt{x-4}\]